Hello
Do you know what the formula to find the area of a shaded region in a circle is?
In this problem I'm working on, the radius is 5m and the shaded region is 75 degrees.
Thank you

Answer by checkley77(12844)

First you need to find the area of the 5m circle.
a=pir^2
a=3.14*5^2
a=3.14*25
a=78.5 m^2 is the area.
Now determine what percentage of the entire circle is contained in 75 degrees.
75/360=.20833 or 20.833 % of the circle.
.20833*78.5=16.35 m^2 is the area within the 75degree slice.